The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu and Dominik J. Bomans

MCnews@astro.uiuc.edu

http://www.astro.uiuc.edu/mcnews/MCnews.html

Number: 4 April 24, 1996

Table of Contents

News and Views..............1 Paper Abstracts..............2

News and Views

“New Views of the Magellanic Clouds”

Topical session in the 188th AAS meeting at Madison, Wisconsin

The 188th AAS meeting at Madison will be held on June 9-13. The special session “New Views of the Magellanic Clouds” is scheduled for Wednesday, June 12. Below is the tentative program for the oral session. Invited talks are 30-35 minutes long, and contributed talks 15 minutes each (including discussion). Poster contributions will be displayed on the same day.

8:30-9:30

Y.-H. Chu  Short Introduction
R.C. Smith  Inv The UM/CTIO Magellanic Cloud Emission-line Survey
J.Wm. Parker  UIT Observations of OB Associations and HII Regions in the MCs

10:15-12:30

S.L. Snowden  Inv Soft X-ray Mosaics of the Large and Small Magellanic Clouds
L. Staveley-Smith Inv An HI Aperture Synthesis Mosaic of the Magellanic Clouds
J.M. Dickey  21-cm Emission-Absorption Studies of the 30 Dor Region
M.L. Kutner  Inv Results of the ESO-SEST Key Programme: “CO in the MCs”
A.D. Bolatto  First Submillimeter CI Observations of the Large Magellanic Cloud

2:00-3:30

D. Zaritsky  Inv A Drift-scan Photometric Survey of the Magellanic Clouds
T.A. Smecker-Hane  The Star-Formation History of the Large Magellanic Cloud
K.H. Cook  Inv Variable Stars from Microlensing Surveys
W.E. Kunkel  The Mass of the Large Magellanic Cloud

3:45-5:30

G.L. White  Inv Radio Continuum Surveys of the Magellanic Clouds
J.R. Dickel  Supernova Remnants in the Magellanic Clouds
N.R. Walborn  Spectral Classification in 30 Dor
A. Caulet  Probing the Multi-Phase ISM of LMC 2
B.J. Weiner  Hα Emission from the Magellanic Stream
M. Irwin  The Proper Motion of the SMC

Hope to see you at Madison,

You-Hua Chu, Dominik Bomans, Bart Wakker

Paper Abstracts

Two nitrogen rich main sequence B-stars in the SMC cluster, NGC 330

D.J. Lennon^{1}, P.L. Dufton^{2}, P.A. Mazzali^{3}, F. Pasian^{3}, G. Marconi^{4}

1 Instititut für Astronomie und Astrophysik der Universität München, Scheinerstrasse 1, D-81679 München, Germany 2 The Department of Pure and Applied Physics, The Queen’s University of Belfast, Belfast BT7 1NN, N. Ireland 3 Osservatorio Astronomico di Trieste, Via G.B. Tiepolo 11, I-34131 Trieste, Italy 4 Osservatorio Astronomico, Viale dell’Osservatorio, I-00040 Monte Porzio Catone (Rome), Italy

High resolution spectra of two narrow lined main-sequence B-type stars in the SMC cluster NGC330 are analysed using LTE methods to obtain their chemical compositions relative to SMC and galactic field B-stars. It is found that these stars, designated A01 and B04, have parameters ({T_{eff}, log g}) of {26 000, 3.8} and {23 000, 3.6} respectively and, surprisingly, nitrogen abundances approximately 0.8 dex higher than other SMC main-sequence B-stars. There is also marginal evidence for moderate deficiencies of both oxygen and silicon consistent with the previously reported low metallicity of this cluster. Since supergiants in the cluster do not exhibit such a large nitrogen enrichment it is argued that A01 and B04 have had their surface chemical compositions modified either as the result of binary mass transfer or due to rotationally induced turbulent diffusion. For both scenarios the absence of a complementary carbon deficiency is surprising. These findings, combined with previous observations of the most lumninous main-sequence stars in the cluster, imply that the upper main-sequence of the cluster is difficult to understand in terms of standard stellar evolution models. Accepted by Astronomy and Astrophysics For preprints, contact djl@usm.uni-muenchen.de

WFPC2 Observations of Star Cluster in the Magellanic Clouds: I. The LMC Globular Cluster Hodge 11

K.J. Mighell^{1} and R.M. Rich^{1}, M. Shara^{2}, S.M. Fall^{2}

1 Dept. of Astronomy, Columbia University, 538 W. 120th Street, New York, NY 10027 2 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218

We present our analysis of Hubble Space Telescope Wide Field Planetary Camera 2 observations in F555W (broadband V) and F450W (broadband B) of the globular cluster Hodge 11 in the Large Magellanic Cloud galaxy. The resulting V vs. (B-V) color-magnitude diagram reaches 2.4 mag below the main-sequence turnoff (which is at V_{T O} = 22.65±0.10 mag or M_{V}^{T O} = 4.00±0.16mag). Comparing the fiducial sequence of Hodge 11 with that of the Galactic globular cluster M92, we conclude that, within the accuracy of our photometry, the age of Hodge 11 is identical to that of M92 with a relative age-difference uncertainty ranging from 10% to 21%. Provided that Hodge 11 has always been a part of the Large Magellanic Cloud and was not stripped from the halo of the Milky Way or absorbed from a cannibalized dwarf spheroidal galaxy, then the oldest stars in the Large Magellanic Clouds and the Milky Way appear to have the same age. Accepted by Astronomical Journal For preprints, contact mighell@figaro.phys.columbia.edu

Extragalactic symbiotic systems IV. The supersoft X-ray source SMC 3

S. Jordan^{1}, W. Schmutz^{2}, B. Wolff^{1}, K. Werner^{1,3}, U. Mürset^{2}

1 Institut für Astronomie und Astrophysik, Universität Kiel, D-24098 Kiel, Germany 2 Institut für Astronomie, ETH Zentrum, CH-8092Zürich, Switzerland 3 Lehrstuhl Astrophysik, Universität Potsdam, Am Neuen Palais 10, D-14469 Potsdam, Germany

We present a consistent model for the UV and supersoft X-ray emission from the symbiotic nova SMC3 (= RX J0048.4–7332). Following the present picture of symbiotic stars, the model consists of radiation from a hot star and an emission nebula excited by that star. The observations were compared to theoretical models in which the hot star’s emission is calculated with the help of hydrostatic and Wolf-Rayet-type non-LTE model atmospheres. Our analysis clearly shows evidence for mass loss rates of several 10^{−6} M_{⊙}/yr. The minimum effective temperature compatible with both the observed UV and X-ray flux is about 260 000 K, which is higher than in any other star analyzed with sophisticated NLTE model atmospheres. Since the hydrostatic surface is hidden by the stellar wind no upper limit for the temperature can be determined. However, we were able to determine the total luminosity of a symbiotic nova with reasonable accuracy (LSMC3 = 10^{4.05±0.05} L_{⊙}). This value is well below the Eddington limit (≈ 50 000 L_{⊙}). In order to reproduce the observed energy distribution a carbon-to-helium ratio > 2 · 10^{−4} — leading to an absorption edge at 0.39 keV — is necessary. Accepted by Astronomy & Astrophysics For preprints, contact jordan@astrophysik.uni-kiel.d400.de

The Dynamics of Superbubbles in the Large Magellanic Cloud

M. S. Oey^{1,2}

1 Steward Observatory, University of Arizona, Tucson, AZ 85721 2 Present address: Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, United Kingdom

Based on the stellar populations observed within this sample of LMC superbubbles, I use a numerical version of the standard, pressure-driven bubble model to investigate the shell dynamics. The results fall into two distinct categories corresponding to a subset of objects for which the observed expansion velocity is too large for the observed shell radius (“high-velocity” superbubbles), and a subset of objects which appears more dynamically consistent with the model (“low-velocity superbubbles”). Both subsets of objects imply an overestimate in the shell growth rate equivalent to an overestimate in input power by up to an order of magnitude. The high-velocity objects exhibit X-ray evidence of supernova activity, suggesting that the dynamical discrepancy is due to acceleration by SNR impacts. Accepted by Astrophysical Journal For preprints, contact oey@ast.cam.ac.uk Postscript version available at http://www.ast.cam.ac.uk/∼oey/oey pubs.html.

A radio continuum study of the Magellanic Clouds Part V: Catalogues of radio sources in the Small Magellanic Cloud

at 1.42, 2.45, 4.75, 4.85 and 8.55 GHz

M. D. Filipović^{1,2}, P. A. Jones^{1}, G. L. White^{1}, R. F. Haynes^{2,1}, U. Klein^{3} and R. Wielebinski^{4}

1 University of Western Sydney Nepean, P.O. Box 10, Kingswood, NSW 2747 Australia 2 Australia Telescope National Facility, CSIRO Radiophysics, P.O. Box 76, Epping, NSW 2121 Australia 3 Radioastronomisches Institut der Universitat Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 4 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany

We present catalogues of radio sources in the Small Magellanic Cloud from observations with the Parkes radio telescope at four frequencies: 1.42, 2.45, 4.75 and 8.55 GHz and an additional catalogue from the Parkes-MIT-NRAO survey at 4.85 GHz. A total of 224 sources have been detected at at least one of these frequencies, 60 of which are reported here for the first time as radio sources. We compare positions and flux densities of these sources with previously published results and find no significant positional displacement or flux discrepancies. Submitted to Astronomy & Astrophysics Suppl. For preprints, contact fica@st.nepean.uws.edu.au